OpenStudy (kellyspeakslouder):

This is a writing portion Find f(x) and g(x) so the function can be expressed as y = f(g(x)). y = Eight divided by x squared. + 4

OpenStudy (kellyspeakslouder):

@jim_thompson5910 my last question, I promise

OpenStudy (kellyspeakslouder):

@DanJS

OpenStudy (kellyspeakslouder):

@zepdrix

satellite73 (satellite73):

\[y=\frac{8}{x^2+4}\]??

OpenStudy (kellyspeakslouder):

yes

satellite73 (satellite73):

you got lots of choices for this one suppose you had to evaluate this at say \(5\) what would you do first if you wanted to compute \[\frac{8}{5^2+4}\]

OpenStudy (kellyspeakslouder):

would this work? f(x) and g(x) are 2(x + 2) and 4x² respectively so that the function can be expressed as y = f(g(x)).

satellite73 (satellite73):

not if it was the one i wrote above,no you need one to be a fraction of some kind

satellite73 (satellite73):

f(x) and g(x) are 2(x + 2) and 4x² then \[f(g(x))=f(4x^2)=2(4x^2+2)\]

OpenStudy (kellyspeakslouder):

oh, sorry. I didn't see what you sent. I was typing this..

satellite73 (satellite73):

go ahead and compute this number, see how you do it

satellite73 (satellite73):

\[\frac{8}{5^2+4}\]

OpenStudy (kellyspeakslouder):

what does compute mean

satellite73 (satellite73):

calculate, find the number, whatever

OpenStudy (kellyspeakslouder):

I keep getting 4.32

satellite73 (satellite73):

seems unlikely since the denominator is larger than the numerator not really interested in the decimal answer, just in the method for computing the number

satellite73 (satellite73):

what is step one in finding \[\frac{8}{5^2+4}\] as in "order of operations" what do you do first ?

OpenStudy (kellyspeakslouder):

the denominator would be 29 right?

satellite73 (satellite73):

yes, so final answer would be \[\frac{8}{29}\] how did you get the 29?

OpenStudy (kellyspeakslouder):

5 squared is 25 plus 4 equals 29

satellite73 (satellite73):

right , step one was square 5, step two add 4, step three put it under 8

satellite73 (satellite73):

so you could say for \[y=\frac{8}{x^2+4}\] since step one is square put \(f(x)=x^2\) and \[g(x)=\frac{8}{x+4}\] but you could do something else

satellite73 (satellite73):

you could combine the first two steps square and add 4, and put \[f(x)=x^2+4,g(x)=\frac{8}{x}\]

satellite73 (satellite73):

either one would work as would others

OpenStudy (kellyspeakslouder):

thank you

satellite73 (satellite73):

yw, hope it is more or less clear

jimthompson5910 (jim_thompson5910):

Another way is to write out \[\Large f(x) = \frac{8}{x^2+4}\] Then highlight the x^2 term in red \[\Large f(x) = \frac{8}{{\color{red}{x^2}}+4}\] We can replace that with some function g(x). So I'm defining g(x) = x^2 \[\Large f(x) = \frac{8}{{\color{red}{x^2}}+4}\] \[\Large f(x) = \frac{8}{{\color{red}{g(x)}}+4}\]

jimthompson5910 (jim_thompson5910):

If we had \[\Large f(x) = \frac{8}{x+4}\] and g(x) = x^2, then evaluating f(g(x)) would give... \[\Large f(x) = \frac{8}{x+4}\] \[\Large f({\color{red}{x}}) = \frac{8}{{\color{red}{x}}+4}\] \[\Large f({\color{red}{g(x)}}) = \frac{8}{{\color{red}{g(x)}}+4}\] \[\Large f({\color{red}{g(x)}}) = \frac{8}{{\color{red}{x^2}}+4}\]